OpenStudy (anonymous):

Divide 5 and one-half divided by one and three-fourths. 3 and one-seventh 9 and five-eighths 5 and two-thirds 3 and one-fourth

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

5 and one-half divided by one and three-fourths. =(5*1/2)/(1*3/4) =(11/2)/(7/4) =11/2*4/7=44/14=22/7= three and one seventh